A car moving along a straight level road has an initial velocity of 16ms^-1. It's deceleration is 2ms^-2. How long does it take for the car to stop?
1
Expert's answer
2013-02-05T09:02:44-0500
a = dv/dt
a - acceleration ( -a - deceleration )
v - velocity
t - time
Therefore:
v = a*t + v0
v0 - initial velocity
In our case:
v = 0 = 16 - 2*t ( car stops)
t = 16/2 = 8 seconds
Answer: 8 seconds.